OpenStudy (anonymous):

Find the perimeter and area of a right triangle if the shortest side is 0.6 m. and the longest side is 1 m. Include correct units with each part of your solution.

OpenStudy (anonymous):

by pythagoras rule the other side is 0.8m so perimeter =1+0.6+0.8=2.4 area = 0.5*0.6*0.8=0.24

OpenStudy (anonymous):

first one in metres, area in metres squared

OpenStudy (anonymous):

?

OpenStudy (anonymous):

2.4m 0.24m^2?

OpenStudy (anonymous):

yh